# Dual formulation of the maximum entropy method applied to analytic continuation of quantum Monte Carlo data

### v1

@@ 1 Introduction
[…]
@@ 3.1 Noisy Linear Model of PIMC Data
[…]
@@
[…]
@@
[…]
@@ 3.2.1 KL-regularized least-squares
where $\|z\|^{2}_{C^{-1}}=z^{T}C^{-1}z$ denotes the scaled 2-norm, $C$ represents the covariance matrix of the error in the observation vector $b$, and regularization parameter $\alpha\geq 0$ is a positive regularization parameter that balances the importance of fitting the observed measurements versus matching the prior $\mu$. The term
@@
is the Shannon–Jaynes (SJ) relative entropy [^2] Relative entropy (Kullback–Leibler Divergence) belongs to the family of $f$-divergences, which are often interpreted as pseudo-metrics for probability measures. between discrete distributions $x$ and $\mu$. Going forward we refer to (10) as the relative entropy. We adopt the standard convention that $00\log 0=$ 0. Note that where $\mu_{i}=0$ then $x_{i}$ must also be zero to render $\operatorname{S_{SJ}}(x\mid\mu)$ finite. The relative entropy encodes the conventional relative entropy term in additional to a linear term
@@
which enforces a soft-constraint on the positive normalization of $x$ with that of $\mu$. Our inversion code allows the user to omit the soft constraint in (11), optimizing instead the conventional relative entropy.
@@ 3.2.3 Dual of KL-regularized least squares
The objective function given in (13), referred to as the primal problem, is nonsmooth, necessitating the use of constrained or nonsmooth optimization algorithms. Such algorithms, however, are typically less efficient and reliable than those designed for smooth unconstrained optimization [43, Section 17.2]. Given the strict convexity of the KL divergence in (15), we can leverage the duality between strict convexity and differentiability [44, Theorem 11.13]. Neglecting $\sum_{j=1}^{n}\mu_{j}-x_{j}$ from (13), the dual of the KL-regularized least-squares problem (13) is the minimization, over $y\in{\mathbb{R}}^{N_{\tau}}$, of the strongly convex and smooth objective function
@@
$d(y,Z\mid\mu):=\frac{\alpha}{2}y^{T}Cy-\langle b,y\rangle+Z\mathop{\mathrm{logexp}}(A^{*}y\mid\mu)-Z\log(Z),$
@@ 3.3.1 Relationship between dual error and primal error
[…]
@@ 3.3.2 Relationship between data perturbations and primal solution
Observe that the solution map $P$ is the proximal map corresponding to the $\mathop{\mathrm{logexp}}$ function (18)  [45, Chapter 6]. Under mild conditions– which are met by $f$– the proximal map is nonexpansive [46, Proposition 23.8]. This means that for any two vectors $b$ and $b^{\prime}$,
@@ 4.1.1 Synthetic correlation samples from zero-temperature $\rho$ meson
For this LQCD benchmark, we consider Asakawa, Hatsuda, and Nakamora’s seminal work which introduced the MEM with Bryan’s algorithm to quarkonia spectral function [10]. In particular, we study the synthetic correlation data generated from the “realistic spectral function” of a charged $\rho$ meson. The charged $\rho$ meson spectral function is parameterized by the relativistic Breit-Wigner function and the relativistic Breit-Wigner’s parameters were fit to reproduce the cross section for $e^{+}e^{-}$ annihilating into hadrons [53]. Spectral functions are by definition $S(k=0,\omega)$ and often denoted as $\rho(\omega)$ in the LQCD community. We consider the spectral function $\rho(\omega)=\omega^{2}x(\omega)$, where $x$ is given
@@
$\sigma_{i}^{2}=\sigma^{2}\frac{\tau_{i}+\epsilon}{\Delta\tau}b_{i}$ (31)
where $\sigma^{2}$ is a noise level we will vary and the perturbation $\epsilon=10^{-6}$ is introduced to ensure $\sigma_{i=0}>0$. Since the noise is proportional to $\tau$ the signal is more precise at small $\tau$, which helps the method to resolve features at large $\omega$. Using this noise we generate $N_{s}=1000$ samples and estimate the data $b$ and the ecovariance matrix $C_{i,j}=\delta_{i,j}\sigma_{i}^{2}/\sqrt{N_{s}}$ from those samples. This differs slightly from Asakawa et al. who generated a single sample and used the true covariance matrix; our method is more reflective of what is done in practice. We normalize the data $b_{i}$ and standard deviation $\sigma_{i}$ by the first entry of the data $b_{0}$. Anecdotally, this makes $\sum_{i}x_{i}<1$. We use a flat prior $\mu=0.257$ as our default model and normalize it by $b_{0}$. We then compute a collection of solutions $x_{\alpha}(\omega)$ over a domain of $\alpha$’s. From this larger $\alpha$ domain, we select a subdomain [$\alpha_{\textrm{min}},\alpha_{\textrm{max}}$] where the posterior distribution– derived by Gull [39]$P(\alpha\mid b,\mu)$ and the maximum of that distribution $P(\alpha^{*}\mid b,\mu)$ are related by the bound $P(\alpha\mid b,\mu)\leq 0.1\times P(\alpha^{*}\mid b,\mu)$. We compute the estimate $\hat{x}$ from the collection of solutions as
@@
where the index $\alpha$ is the $\alpha$ values within [$\alpha_{\textrm{min}},\alpha_{\textrm{max}}$] and $x_{\alpha}$ is the corresponding solution. The posterior weighting function is
@@
Here $\lambda_{k}$ is the $k^{\mathrm{th}}$ eigenvalue of the real symmetric matrix $\Lambda_{j,j^{\prime}}(\alpha)=\sqrt{x_{j}(\alpha)}\frac{\partial^{2}L}{\partial x_{j}(\alpha)\partial x_{j^{\prime}}(\alpha)}\sqrt{x_{j^{\prime}}(\alpha)}$. In the typical fashion, we estimate the error as ±$\pm 2\sqrt{\widehat{(x^{2})}-(\hat{x})^{2}}$.
@@
Figure 1: Plots of the parameterized $\rho$-meson spectral function (29), the MEM estimates, and the flat Bayesian prior. Plots keep number of correlator points constant (i.e., $N_{\tau}=30$) and vary the noise level $\sigma$. From left to right the noise increases as $\sigma^{2}=10^{-4},10^{-3},10^{-2}$. In the leftmost plot ($\sigma^{2}=10^{-4}$), the estimates and error bands are identical except near $\omega=0$.
@@
Figure 2: In rows 1 and 2 respectively, we present plots of the solutions proposed by the dual Newton algorithm and Bryan’s algorithm at different regularization weights $\alpha$. All inversions use synthetic $\rho$-meson data generated from $N_{\tau}=30$, but the noise parameter increases from left to right, i.e., $\sigma^{2}=10^{-3}$ (left column), $\sigma^{2}=10^{-2}$ (middle column), $\sigma^{2}=10^{-1}$. (right column). In the bottom row we present the posterior weighting function used to combine the various solutions. The red and blue color bands indicate the $\alpha$ domain $[\alpha_{\mathbf{min}},\alpha_{\mathbf{max}}]$ kept by the MEM with dual Newton optimizer and MEM with Bryan’s optimizer respectively; see (32) for details. In all cases the posterior weighting function associated with the dual Newton method has a sharper peak at smaller $\alpha$ value than Bryan’s primal approach. Comparing posterior weighting function across columns, as noise increases the peak of the posterior weighting function for Bryan’s primal approach moves rightward. In comparison, the peak of the posterior weighting function for the dual Newton moves less and in the leftward direction.
@@ 4.2.1 Synthetic Correlation Data from UEG Dynamic Structure Factor
The uniform electron gas (UEG) [41, 55, 56] is one of the most well-studied systems in statistical physics and quantum chemistry, making it very suitable for assessing an algorithm’s performance and uncertainty quantification. A DSF can be produced from a susceptibility $\chi$ via the fluctuation dissipation theorem [41],
$S(k,\omega)=\frac{2\hbar}{n}\frac{\text{Im}\chi(k,\omega)}{1-e^{-\beta\hbar\omega}},$ (33)
@@
In this test problem, we generate synthetic samples by multiplying the Laplace kernel with the mean field corrected completed Mermin DSF model. We consider the UEG at density $n_{e}=1.611\times 10^{21}\,\text{cc}^{-1}$ which corresponds to a Wigner-seitz radius of $10$ Bohr. The temperature is $T=0.5$ eV, equivalently $\beta=54.301$ Hartree, which corresponds to a quantum degeneracy parameter of $\Theta=T/E_{F}=1$; these conditions match the conditions of the PIMC UEG simulation, whose correlation estimates will be considered next. For this $n$ and $T$, the plasma frequency is $\omega_{p,e}=\sqrt{4\pi ne^{2}/m_{e}}=0.05476$ Hartree and Debye wavenumber is $k_{D,e}=\sqrt{4\pi ne^{2}/T}=0.40362\,\text{Bohr}^{-1}$. The collision frequency $\nu$ is set equal to the plasmon frequency $\omega_{p,e}$. We discretize our frequency as $\omega_{j}=j\Delta\omega$ where $j\in\{1,\ldots,N_{\omega}\}$, $N_{\omega}=1250$, $\Delta\omega=0.0011$ Hartree. Additionally, we discretize our imaginary time as $\tau_{i}=i\Delta\tau$ where $i\in\{0,\ldots,N_{\tau}-1\}$, $N_{\tau}=201$, $\Delta\tau=0.271505$ Hartree. We introduce noise different from (31); we chose a standard deviation of
@@
We first consider the DSF estimate for a small wavenumber $k=0.3990\,\text{Bohr}^{-1}$ and noise levels $\sigma=10^{-3},10^{-2},10^{-1}$. As the noise goes to zero both the MEM with Bryan’s algorithm and the MEM with dual Newton converge to the desired $S(k,\omega)$ and have comparable performance, which is unlike the $\rho$-meson investigation. We interpret this as a result of the smooth nature of the desired signal.
Figure 4: Plots of the estimated dynamic structure factor (units of $\text{eV}^{-1}$) for the synthetic UEG example with associated posterior weighting function. From left to right, the noise level $\sigma$ increases as $\sigma=10^{-3},10^{-2},10^{-1}$. The top row contains the plots of the DSF. The completed Mermin (CM) model is the the input spectrum and desired recovery, defined in (35) with mean field correction (36) included. The RPA UEG is the prior model, defined in (34) with mean field correction (36) included. The MEM estimates, which overlap strongly, are in either red or blue. On the bottom row the Bayesian posterior is plotted as a function of $\alpha$. We logarithmically sample alpha in $\alpha\in[10^{0}-10^{5}]$. The red and blue vertical color bands indicates the $\alpha$ domain is selected for the MEM averaging procedure; see (32) for details on how the $\alpha$ domain was selected.
We conduct inversions across $k$ to estimate the entire $S(k,\omega)$ and observe the dispersion relation across $k$’s; we use noise level $\sigma=0.001$ to generate synthetic data. We find that the plasmon peak is broader and flatter at large $k$ in the RPA approximation than in the completed Mermin approximation. However, both dual and primal MEM estimates preserve the magnitude of the peak at larger $k$, resembling the behavior of the completed Mermin model.
Figure 5: Heatmaps of the dynamic structure factor $S(k,\omega)$. completed Mermin UEG is the input spectrum and desired recovery, defined in (35) with mean field correction (36) included. RPA is the prior model, defined in (34) with mean field correction (36) included. For each $k$ value the location of the DSF peak is indicated with a black dot; together they indicate the dispersion relation.
@@ 4.2.2 Authentic Correlation Data from UEG PIMC simulation
[…]
@@
We present plots of the reconstructed $S(k,\omega)$ at a small, middle and large wavenumber $k=0.1702$, $0.3990$, $0.5513$ $\text{Bohr}^{-1}$. These plots show excellent agreement with the stochastic method at large $k$, but poorer agreement at small $k$. Anecdotally, we found that if the default model is sharply peaked, then the solutions deviate less from the default model. Furthermore, both MEM approaches produce solutions that have less ringing than the stochastic model; ringing at large $\omega$ is a common concern for the BRM entropic approach [7]. Plots are given in Figure 6 top row.
@@
We also present the normalized Bayesian posterior used to combine the $\alpha$-solutions into an estimate. Since the error in the correlation function is small, we do not expect that Bryan’s Bayesian posterior will differ from the dual Newton’s Bayesian posterior. This is reflected in Figure 6 bottom row.
@@
Figure 6: Top: From left to right we increase the wavenumber $k$ considered in our DSF. The curve labeled “stochastic” is Tobias et al.’s estimate, not the exact solution [29]. The RPA UEG is the prior model, defined in (34), corrected with the random-phase approximation (36). Bottom: Plot of the posterior weighting function associated with the above estimates. We collect solutions for $\alpha\in[10^{-3}-10^{3}]$. The red and blue color bands indicate the $\alpha$ domain kept by the MEM with dual Newton optimizer and MEM with Bryan’s optimizer respectively; see (32) for details on how the $\alpha$ domain was selected.
@@
We plot the correlation function that would be produced from the reconstructed DSFs and compare to the PIMC UEG data. This helps us compare to the stochastic sampling method which accepted and averaged over DSF’s based on the proposed DSF’s ability to “reproduce the [correlation function] within the Monte Carlo error bars”. We see that both MEM with the dual Newton and Bryan’s optimization produce a correlation functions that are indistinguishable from the data. The associated reduced chi-squared value $\chi^{2}/N_{\tau}=(Ax-b)C^{-1}(Ax-b)/N_{\tau}$ reflects this. The chi-square values also inform us that the MEM estimates match the data more closely than the stochastic procedure. Plots of the correlation functions are given in Figure 7.
Figure 7: These plots present the signal associated each of the estimates presented in the top row of Figure 6. We include the observed PIMC UEG signal for comparison, but the signal associated to the estimates is visually indistinguishable from the observed PIMC UEG signal in all cases except the for RPA DSF.
@@
We also investigate the full DSF $S(k,\omega)$ estimate and find that both Bryan’s algorithm and the dual Newton algorithm produces DSF’s that contain a dip in the dispersion relation occurring for $k\in[0.25,0.45]\,\text{Bohr}^{-1}$, which is known as a roton feature [63]. Comparing to the stochastic approach the MEM approaches do not produce as smooth of a dispersion relation. We interpret this as a result of the data indicating there is a roton feature, but the prior model (i.e., RPA) not having this feature and then some compromise is reached. This serves as a reminder that all forms of entropic regularization produce a unique solution that is biased towards the default model. Two possible ways to address this are using improved default models such as the static approximation in the case of the warm dense UEG [64] or to conduct the entire $S(k,\omega)$ inversion at once rather than divided into individual $k$ inversions. Plots of the full DSF are given in Figure 8.
@@
Figure 8: Heatmaps of the dynamic structure factor $S(k,\omega)$. The $S(k,\omega)$ estimate from [61] is labeled “stochastic sampling”. The prior model, indicated “RPA” is defined in (34) with mean field correction (36) included. The estimate obtained by MEM with Bryan’s algorithm is indicated by “Bryan” and the estimate obtained by MEM with the dual Newton optimization is indicated by “dual Newton”. For each $k$ value the location of the DSF peak is indicated with a black dot; together they indicate the dispersion relation. Dashed white lines have been placed at $k=0.170153$, $0.399044$, and $0.551358\,\text{Bohr}^{-1}$. These lines indicate the $k$ values plotted in Figures 6 and 7

### v2

@@ 1 Introduction
$F(\tau)=\frac{1}{2\pi}\int_{-\infty}^{+\infty}A(\tau,\omega)S(\omega)\,d\omega,$ (1)
@@ 3.1 Noisy Linear Model of PIMC Data
$F(k,\tau)\equiv\frac{1}{2\pi}\int_{-\infty}^{+\infty}\exp(-\omega\tau)S(k,\omega)d\omega+\eta(k,\tau).$ (3)
@@
$F(k,\tau)\equiv\frac{1}{2\pi}\int_{0}^{\infty}\left(\exp(-\omega\tau)+\exp(-\omega(\beta-\tau))\right)S(k,\omega)d\omega+\eta.$ (4)
@@
$F(k,\tau=0)=\frac{1}{2\pi}\int_{0}^{\infty}\left(1+\exp(-\omega\beta)\right)S(k,\omega)d\omega.$ (7)
@@ 3.2.1 KL-regularized least-squares
where $\|z\|^{2}_{C^{-1}}=z^{T}C^{-1}z$ denotes the scaled 2-norm, $C$ represents the covariance matrix of the error in the observation vector $b$, and regularization parameter $\alpha\geq 0$ is a positive regularization parameter that balances the importance of fitting the observed measurements versus matching the Bayesian prior $\mu$. The term
@@
is the Shannon–Jaynes (SJ) relative entropy [^2] Relative entropy (Kullback–Leibler Divergence) belongs to the family of $f$-divergences, which are often interpreted as pseudo-metrics for probability measures. between discrete distributions $x$ and $\mu$. Going forward we refer to (10) as the relative entropy. We adopt the standard convention that $0\log 0=0$. Note that where $\mu_{i}=0$ then $x_{i}$ must also be zero to render $\operatorname{S_{SJ}}(x\mid\mu)$ finite. The relative entropy encodes the conventional relative entropy term in additional to a linear term
@@
which enforces a soft-constraint on the positive normalization of $x$ with that of $\mu$. As an option, our inversion code allows the user to omit the soft constraint in (11), optimizing instead the conventional relative entropy. In this work, we include the soft constraint to make direct comparisons.
@@ 3.2.3 Dual of KL-regularized least squares
The objective function given in (13), referred to as the primal problem, is nonsmooth, necessitating the use of constrained or nonsmooth optimization algorithms. Such algorithms, however, are typically less efficient and reliable than those designed for smooth unconstrained optimization [43, Section 17.2]. Given the strict convexity of the KL divergence in (15), we can leverage the duality between strict convexity and differentiability [44, Theorem 11.13]. The dual of the KL-regularized least-squares problem (13) is the minimization, over $y\in{\mathbb{R}}^{N_{\tau}}$, of the strongly convex and smooth objective function
@@
$d(y,Z\mid\mu):=\frac{\alpha}{2}y^{T}Cy-\langle b,y\rangle+Z\mathop{\mathrm{logexp}}(A^{*}y-c\mid\mu)-Z\log(Z),$ (18)
@@ 3.3.1 Relationship between dual error and primal error
where $\|\cdot\|$ denotes the 2-norm, and hence $\|A_{C}\|$ is the largest singular value of $A_{C}$. Equation (21) shows that any error in the inferred primal solution is bounded by a constant multiple of the error in the dual solution. An immediate application of this bound concerns the sequence of dual iterates $y^{k}$ generated by the Newton-Krylov method described in A. Specifically, the error in corresponding primal iterate $x(y^{k})$ is bounded by the error in $y^{k}$ scaled by the norm of the operator $A_{C}$:
@@ 3.3.2 Relationship between data perturbations and primal solution
Observe that the solution map $P$ is the proximal map corresponding to the $\mathop{\mathrm{logexp}}$ function (19)  [45, Chapter 6]. Under mild conditions– which are met by ff– the proximal map is nonexpansive [46, Proposition 23.8]. This means that for any two vectors $b$ and $b^{\prime}$,
@@ 4.1.1 Synthetic correlation samples from zero-temperature $\rho$ meson
For this LQCD benchmark, we consider Asakawa, Nakahara, and Hatsuda’s seminal work which introduced the MEM with Bryan’s algorithm to quarkonia spectral function [10]. In particular, we study the synthetic correlation data generated from the “realistic spectral function” of a charged $\rho$ meson. The charged $\rho$ meson spectral function is parameterized by the relativistic Breit-Wigner function and the relativistic Breit-Wigner’s parameters were fit to reproduce the cross section for $e^{+}e^{-}$ annihilating into hadrons [53]. Spectral functions are by definition $S(k=0,\omega)$ and often denoted as $\rho(\omega)$ in the LQCD community. We consider the spectral function $\rho(\omega)=\omega^{2}x(\omega)$, where $x$ is given
@@
The empirical values of the constants are given by $m_{\rho}=0.77$ GeV, $m_{\pi}=0.14$ GeV, $g_{\rho\pi\pi}=5.45$, $\omega_{0}=1.3$ GeV, $\delta=.02$ GeV. We multiply $x(\omega)$ with kernel $A_{i,j}=e^{\tau_{i}\omega_{j}}\omega_{j}^{2}$ to produce our noiseless synthetic correlation data $b$. We discretize using $\omega_{j}=j\Delta\omega$ where $j\in\{1,\ldots,N_{\omega}\}$, $N_{\omega}=600$, and $\Delta\omega=0.01$ GeV as well as $\tau_{i}=i\Delta\tau$ where $i\in\{0,\ldots,N_{\tau}-1\}$, $N_{\tau}=30$, and $\Delta\tau=0.431\,\text{GeV}^{-1}$. We introduce noise $\eta(\tau_{i})\sim\mathcal{N}(0,\sigma_{i}^{2})$ to the clean signal to construct synthetic samples. Following Asakawa et al., the variance is given by
$\sigma_{i}^{2}=\sigma^{2}\frac{\tau_{i}+\epsilon}{\Delta\tau}b_{i}$ (32)
@@
where the index $\alpha$ is the $\alpha$ values within $\alpha_{\textrm{min}},\alpha_{\textrm{max}}$] and $x_{\alpha}$ is the corresponding solution. The posterior weighting function is
@@
Here $\lambda_{k}$ is the $k^{\mathrm{th}}$ eigenvalue of the real symmetric matrix $\Lambda_{j,j^{\prime}}(\alpha)=\sqrt{x_{j}(\alpha)}\frac{\partial^{2}L}{\partial x_{j}(\alpha)\partial x_{j^{\prime}}(\alpha)}\sqrt{x_{j^{\prime}}(\alpha)}$. In the typical fashion, we estimate the error as $\pm 2\sqrt{\widehat{(x^{2})}-(\hat{x})^{2}}$.
@@
Figure 1: Plots of the parameterized $\rho$-meson spectral function (30), the flat Bayesian prior, and the MEM estimates obtained using either Bryan’s optimization algorithm (blue) or our dual Newton optimization algorithm (red). Plots keep number of correlator points constant (i.e., $N_{\tau}=30$) and vary the noise level $\sigma$. From left to right the noise increases as $\sigma^{2}=10^{-4},10^{-3},10^{-2}$. In the leftmost plot ($\sigma^{2}=10^{-4}$), the estimates and error bands are identical except near $\omega=0$. In the noiseless limit $\sigma^{2}=10^{-4}$ the different optimization algorithms produce similar curves, but when noise is large, MEM with Bryan’s algorithm over-regularizes the solution.
@@
Figure 2: In rows 1 and 2 respectively, we present plots of the solutions proposed by the dual Newton algorithm and Bryan’s algorithm at different regularization weights $\alpha$. All inversions use synthetic $\rho$-meson data generated from $N_{\tau}=30$ and the noise parameter increases from left to right, i.e., $\sigma^{2}=10^{-4}$ (left column), $\sigma^{2}=10^{-3}$ (middle column), $\sigma^{2}=10^{-2}$. (right column). In the bottom row we present the posterior weighting function used to combine the various solutions. The red and blue color bands indicate the $\alpha$ domain $\alpha_{\mathbf{min}},\alpha_{\mathbf{max}}]$ kept by the MEM with dual Newton optimizer and MEM with Bryan’s optimizer respectively; see (33) for details. In all cases the posterior weighting function associated with the dual Newton method has a sharper peak at smaller $\alpha$ value than Bryan’s primal approach. Comparing posterior weighting function across columns, as noise increases the peak of the posterior weighting function for Bryan’s primal approach moves rightward. In comparison, the posterior weighting function for the dual Newton changes little.
@@ 4.2.1 Synthetic Correlation Data from UEG Dynamic Structure Factor
The uniform electron gas (UEG) [41, 55, 56] is one of the most well-studied systems in statistical physics and quantum chemistry, making it very suitable for assessing an algorithm’s performance and uncertainty quantification. A DSF can be produced from a susceptibility $\chi$ via the fluctuation dissipation theorem [41], in atomic units, as
$S(k,\omega)=-\frac{2}{n}\frac{\text{Im}\chi(k,\omega)}{1-e^{-\beta\hbar\omega}},$ (34)
@@
In this test problem, we generate synthetic samples by multiplying the Laplace kernel with the mean field corrected completed Mermin DSF model. We consider the UEG at density $n_{e}=1.611\times 10^{21}\,\text{cc}^{-1}$ which corresponds to a Wigner-seitz radius $r_{s}=10$ Bohr. The temperature is $k_{B}T=0.5$ eV, equivalently $\beta=1/k_{B}T=54.301$ Hartree, which corresponds to a quantum degeneracy parameter of $\Theta=k_{B}T/E_{F}=1$; these conditions match the conditions of the PIMC UEG simulation, whose correlation estimates will be considered next. For this system the electron plasma frequency is $\omega_{p,e}=\sqrt{3/r_{s}^{3}}=0.05476\,\text{Hartree}$ and Fermi wavenumber is $k_{F,e}=(9\pi/4)^{1/3}/r_{s}=0.1919\,\text{Bohr}^{-1}$. The collision frequency $\nu$ is set equal to the electron plasmon frequency $\omega_{p,e}$. We discretize our frequency as $\omega_{j}=j\Delta\omega$ where $j\in\{1,\ldots,N_{\omega}\}$, $N_{\omega}=1250$, $\Delta\omega=0.0011$ Hartree. Additionally, we discretize our imaginary time as $\tau_{i}=i\Delta\tau$ where $i\in\{0,\ldots,N_{\tau}-1\}$, $N_{\tau}=201$, $\Delta\tau=0.271505$ Hartree. We introduce noise different from (32); we chose a standard deviation of
@@
We first consider the DSF estimate for a small wavenumber $k=0.3990\,\text{Bohr}^{-1}$ and noise levels $\sigma=10^{-2},10^{-1},10^{0}$ (see Figure 4). In the limit of small noise, both the MEM with Bryan’s algorithm and the MEM with dual Newton converge to the desired completed Mermin $S(k,\omega)$. However, as noise gets large, we see again that Bryan’s algorithm breaks down, over-regularizing the solution by selecting weighting the larger $\alpha$ values. This is the same behavior seen in for $\rho$-meson investigation.
Figure 4: Plots of the estimated dynamic structure factor for the synthetic UEG example with associated posterior weighting function. From left to right, the noise level $\sigma$ increases as $\sigma=10^{-2},10^{-1},10^{0}$. The top row contains the plots of the DSF. The completed Mermin (CM) model is the the input spectrum and desired recovery, defined in (36) with mean field correction (37) included. The RPA UEG is the Bayesian prior, defined in (35) with mean field correction (37) included. The MEM estimates are in either red or blue. On the bottom row the Bayesian posterior is plotted as a function of $\alpha$. We logarithmically sample alpha in $\alpha\in[10^{0}-10^{5}]$. The red and blue vertical color bands indicates the $\alpha$ domain is selected for either MEM averaging procedure; see (33) for details on how the $\alpha$ domain was selected. Comparing posterior weighting functions, as noise increases the peak of the posterior weighting function for Bryan’s primal approach moves rightward. In comparison, the posterior weighting function for the dual Newton changes little.
@@ 4.2.2 Authentic Correlation Data from UEG PIMC simulation
The authentic WDM UEG PIMC correlation data we study arises from Dornheim and Vorberger’s work [61]. They employed path integral Monte Carlo (PIMC) methods to produce Euclidean time correlation functions of the uniform electron gas at WDM conditions (T=0.5T=0.5eV and $r_{s}=10$). This PIMC UEG data was analyzed by Dornheim et al. [29], who used a stochastic sampling algorithm to infer the DSF. Their stochastic sampling algorithm is a forward problem approach that relies on sampling values for a parameterized form of the susceptibility; the parameterized form is determined via sum rules and additional constraints. By comparison, our inverse method has no similar guarantee of satisfying sum rules, but will work in systems where a parameterized form of the susceptibility is not known and hence has a broader range of applicability. In this section, we include the stochastic sampling results in the slot where the completed Mermin was in synthetic results because the general shape of the DSF is physical even if the stochastic estimate is not the true solution.
@@
We present plots of the reconstructed $S(k,\omega)$ at a small, middle and large wavenumber $k=0.1702$, $0.3990$, $0.5513$ $\text{Bohr}^{-1}$. These plots show excellent agreement with the stochastic method at large $k$, but poorer agreement at small $k$. Anecdotally, we found that if the Bayesian prior is sharply peaked, then the solutions deviate less from the prior. Furthermore, both MEM approaches produce solutions that have less ringing than the stochastic model; ringing at large $\omega$ is a common concern for the BRM entropic approach [7]. Plots are given in Figure 5 top row.
@@
We also present the normalized Bayesian posterior used to combine the $\alpha$-solutions into an estimate. Since the error in the correlation function is small, approximately $10^{-3}$, we do not expect that Bryan’s Bayesian posterior will differ from the dual Newton’s Bayesian posterior. This is reflected in Figure 5 bottom row and explains why the MEM with Bryan and the MEM with dual Newton estimates are indistinguishable from the PIMC data..
@@
Figure 5: Top: From left to right we vary the wavenumber $k$ considered in our DSF, for all $k$ values, the number of correlator points is fixed at $201$ and the relative noise is $\approx 10^{-3}$. The curve labeled “stochastic” is Tobias et al.’s estimate [29], not the exact solution. The RPA UEG is the Bayesian prior, defined in (35), with the mean-field correction (37) included. Bottom: Plot of the posterior weighting function associated with the above estimates. We collect solutions for $\alpha\in[10^{-3}-10^{3}]$. The red and blue color bands indicate the $\alpha$ domain kept by the MEM with dual Newton optimizer and MEM with Bryan’s optimizer respectively; see (33) for details on how the $\alpha$ domain was selected. We find strong overlap between the Bryan and dual Newton approach is readily explained by the small relative noise in the data.
@@
We plot the correlation function that would be produced from the reconstructed DSFs and compare to the PIMC UEG data. This helps us compare to the stochastic sampling method which accepted and averaged over DSF’s based on the proposed DSF’s ability to “reproduce the [correlation function] within the Monte Carlo error bars”. We see that both MEM with the dual Newton and Bryan’s optimization produce correlation functions that are indistinguishable from the data. The associated reduced chi-squared value $\chi^{2}/N_{\tau}=(Ax-b)C^{-1}(Ax-b)/N_{\tau}$ reflects this. The chi-square values also inform us that the MEM estimates match the data more closely than the stochastic procedure. Plots of the correlation functions are given in Figure 6.
Figure 6: Plots of the periodic Laplace transformed DSF estimates presented in the top row of Figure 5 (transform defined in  (4)). We include the imaginary time correlation function (ITCF) data obtained via PIMC UEG data for comparison. The periodic Laplace transform of the MEM DSF estimates are visually indistinguishable from the observed ITCF data, while the periodic Laplace transform RPA DSF shows clear differences. This indicates that the solution is not so heavily regularized that the entropic prior dominates the cost function.
@@
We also investigate the full DSF $S(k,\omega)$ estimate and find that both Bryan’s algorithm and the dual Newton algorithm produces DSF’s that contain a dip in the dispersion relation occurring for $k\in[0.25,0.45]\,\text{Bohr}^{-1}$, which is known as a roton feature [62]. Comparing to the stochastic approach the MEM approaches do not produce as smooth of a dispersion relation. We interpret this as a result of the data indicating there is a roton feature, but the Bayesian prior (i.e., RPA) not having this feature and then some compromise is reached. This serves as a reminder that all forms of entropic regularization produce a unique solution that is biased towards the Bayesian prior. Two possible ways to address this are using improved default models such as the static approximation in the case of the warm dense UEG [63] or to conduct the entire $S(k,\omega)$ inversion at once rather than divided into individual $k$ inversions. Plots of the full DSF are given in Figure 7.
@@
Figure 7: Heatmaps of the dynamic structure factor $S(k,\omega)$. The $S(k,\omega)$ estimate from [61] is labeled stochastic sampling. The Bayesian prior, indicated RPA is defined in (35) with mean field correction (37) included. The estimates obtained by MEM with Bryan’s algorithm and the estimates obtained by MEM with the dual Newton optimization are also included. Dashed white lines have been placed at $k=0.170153$, $0.399044$, and $0.551358\,\text{Bohr}^{-1}$. These lines indicate the $k$ values plotted in Figures 5 and 6. For each $k$ value the location of the DSF peak is indicated with a black dot; together they indicate the dispersion relation. We see that the MEM DSF estimates have a qualitatively different dispersion relation than the RPA Bayesian priod, which better resembles the stochastic estimate.